# Transversal Hamilton cycles in digraph collections

### v1

@@ 1 Introduction
A transversal $X$ over a collection =$\mathcal{F}=\{F_{1},\ldots,F_{m}\}$ of objects represents an object that intersects every $F_{i}$. Here, the object can be any mathematical object, such as sets, spaces, set systems, matroids, and so forth. Several classical theorems, which guarantee the existence of certain objects, have been restated in the context of transversals, illustrating the feasibility or infeasibility of obtaining such an object as a transversal over a certain collection. To name a few, transversals are extensively studied for Carathéodory’s theorem [4], Helly’s theorem [22], Erdős-Ko-Rado theorem [2], Rota’s basis conjecture [20, 24], and others.
@@
For a given collection $\mathcal{G}=\{G_{1},\ldots,G_{m}\}$ of graphs/hypergraphs/digraphs with the same vertex set $V$, an $m$-edge graph/hypergraph/digraph $H$ with vertices in $V$ is transversal in $\mathcal{G}$ if there exists a bijection $\phi:E(H)\to[m]$ such that $e\in E(G_{\phi(e)})$ for all $e\in E(H).$
By interpreting each $G_{i}$ as the set of edges colored with color $i$, the function $\phi$ is often referred to as a coloring. We say that the coloring $\phi$ is rainbow if it is injective, and $H$ is a rainbow subgraph if there exists a rainbow coloring on it. Many classical results in extremal graph theory have been extended to such transversal settings, exhibiting interesting phenomena. The central question in studying transversal generalizations can be stated as follows.
@@
Let $H$ be a graph/hypergraph/digraph with $m$ edges, $\mathcal{G}=\{G_{1},G_{2},\dots,G_{m}\}$ be a collection of graphs/hypergraphs/digraphs on a common vertes set of size $n$. Which property $\mathcal{P}$ imposed on $\mathcal{G}$ will guarantee the existence of a transversal copy of $H$?
Since every object in the family $\mathcal{G}$ can be the same, a necessary condition for a positive answer to the above question is that every $n$-vertex graph/hypergraph/directed graph satisfying the property $\mathcal{P}$ contains a copy of $H$. However, this alone is not always sufficient. For example, Aharoni, DeVos, de la Maza, Montejano, and Šámal [1] proved that if $\mathcal{G}=\{G_{1},G_{2},G_{3}\}$ is a collection of graphs on a common vertex set of size $n$ and $|E(G_{i})|>(\frac{26-2\sqrt{7}}{81})n^{2}$ for $i\in[3]$, then $\mathcal{G}$ contains a rainbow triangle. Moreover, the constant $\frac{26-2\sqrt{7}}{81}$ is optimal, while Mantel’s theorem states that any $n$-vertex graph with more than $\lfloor\frac{n^{2}}{4}\rfloor$ edges must contain a triangle.
@@
A natural next step is to study rainbow analogues in digraph collections. Cheng, Han, Wang and Wang [8] studied the transversal version of Hajnal-Szemerédi theorem in digraph collections, which asymptotically generalized the corresponding results by Czygrinow-DeBiasio-Kierstead-Molla [15] and Treglown [30]. It is well-known that every tournament contains a Hamilton path, and every strongly connected tournament contains a Hamilton cycle. Chakraborti, Kim, Lee and Seo [6] proved the existence of a transversal directed Hamilton path and cycle in a family of tournaments. They [7] also proved that in a digraph collection consisting of tournaments, there exist transversal Hamilton cycles of all possible orientations possibly except the consistently oriented one. These results generalized the classical theorem of Thomason [28]. Babiński, Grzesik and Prorok [3] studied the transversal Mantel’s theorem in digraph and oriented graph collections. Gerbner, Grzesik, Palmer and Prorok [18] determined the minimum number of edges in each color which guarantees the existence any rainbow directed star.
@@ 1.1 Main result
In this paper, we study the transversal problem in digraph collections. Let $D$ be a digraph with no loops and at most one edge in each direction between every pair of vertices. Denote the set of vertices of $D$ by $V(D)$. Let $\delta^{+}(D)$ and $\delta^{-}(D)$ be the minimum out-degree and minimum in-degree of $D$, respectively. Let $\delta^{0}(D)=\min\{\delta^{+}(D),{\delta^{-}}(D)\}$ be the minimum semi-degree of $D.$ Let $\mathcal{D}=\{D_{1},D_{2},\dots,D_{m}\}$ be a collection of not necessarily distinct digraphs with common vertex set $V$. Define $\delta^{0}(\mathcal{D}):=\min\{\delta^{0}(D_{i}):i\in[m]\}$.
In 1960, Ghouila-Houri [19] generalized the Dirac’s theorem to digraphs.
@@
Every $n$-vertex digraph $D$ with $\delta^{0}(D)\geq\frac{n}{2}$ contains a directed Hamilton cycle.
Chakraborti, Kim, Lee, Seo [6] suggested to consider a transversal version of Ghouila-Houri’s theorem would be interesting. Our main result solves this problem.
@@
In an undirected graph, one may replace each undirected edge by two arcs with opposite directions. Thus, for sufficiently large $n$, Theorem 1.2 is a direct consequence of Theorem 1.4. Furthermore, by letting $D_{1}=D_{2}=\cdots=D_{n}$ in Theorem 1.4, we get Theorem 1.3 immediately when $n$ is large.
@@ 1.2 Notation and Organization
We use $H\subseteq D$ to denote that $H$ is a sub-digraph of $D$. For $S\subseteq V(D)$, we write $D[S]$ for the sub-digraph of $D$ induced by $S$ and $D-S$ for $D[V(D)\setminus S]$. For two not necessarily disjoint vertex subsets $A,B\subseteq V(D)$, let $D[A,B]$ be the sub-digraph of $D$ with vertex set $A\cup B$ and edge set $\{xy\in E(D):x\in A,y\in B\}$ and let $e_{D}(A,B):=|E(D[A,B])|$. Denote $D^{\pm}[A,B]=D[A,B]\cup D[B,A]$. When $A=B$, we often abbreviate $D[A,B]$ and $e_{D}(A,B)$ as $D[A]$ and $e_{D}(A)$ respectively.
For a digraph $D$ and $v\in V(D)$, we denote the set of out-neighbours and in-neighbours of $v$ by $N_{D}^{+}(v)$ and $N_{D}^{-}(v)$, respectively, and we write $d_{D}^{+}(v):=|N_{D}^{+}(v)|$ and $d_{D}^{-}(v):=|N_{D}^{-}(v)|$ for the out- and indegree of $v$ in $D$, respectively. For $S\subseteq V(D)$, we write $d_{D}^{+}(v,S):=|N_{D}^{+}(v)\cap S|$ and $d_{D}^{-}(v,S):=|N_{D}^{-}(v)\cap S|$. We often omit subscripts when they are clear from the context.
Let $\mathcal{D}=\{D_{i}:i\in[m]\}$ be a collection of digraphs on a common vertex set $V$ and $H$ be a rainbow digraph inside $\mathcal{D}$. For a vertex $v\in V$, let $N_{i}^{+}(v):=N_{D_{i}}^{+}(v)$, $N_{i}^{-}(v):=N_{D_{i}}^{-}(v)$, $d_{i}^{+}(v):=N_{i}^{+}(v)$ and $d_{i}^{-}(v):=N_{i}^{-}(v)$. For any two vertex subsets $X,\,Y\subseteq V$, define $\mathcal{D}[X]:=\{D_{i}[X]:i\in[m]\}$, $\mathcal{D}[X,Y]:=\{D_{i}[X,Y]:i\in[m]\}$ and $\mathcal{D}^{\pm}[X,Y]:=\{D_{i}^{\pm}[X,Y]:i\in[m]\}$. If $E(D)=\emptyset$, then we simply write $D=\emptyset$. Similarly, if there is no edge in $D_{i}$ for all $i\in[m]$, then we write $\mathcal{D}=\emptyset$. Denote by ${\rm col}(H)$ the set of colors appearing in $H$. Two rainbow graphs $H_{1}$ and $H_{2}$ are said to be disjoint if $V(H_{1})\cap V(H_{2})=\emptyset$ and ${\rm col}(H_{1})\cap{\rm col}(H_{2})=\emptyset$.
@@
Organization. We conclude this section with a sketch of the proof of Theorem 1.4. In section 2, we introduce two key techniques that we will need later, i.e., the regularity lemma in digraph collections and a stability result for transversal perfect matchings in bipartite graph collections. For readability, we postpone its proofs to the appendix. The proof of Theorem 1.4 will be divided into “stable case” and “extremal case”. Section 3 introduces notions of extremality and stability for digraphs and digraph collections; we prove several results about absorption that will be used in the proof of the stable case. In Section 5, we show that there is always a transversal directed Hamilton cycle in stable case. Section 6 deals with the extremal case. We finish our paper with some concluding remarks in the final section.
@@ 1.3 Discussion of the strategy
Joos and Kim [21] proved the transversal version of Dirac’s theorem by introducing the auxiliary digraph technique, along with some rotations. However, this technique is not feasible in digraph collections because we cannot extend an existing directed cycle into a longer cycle by using the same rotation. In [10], the first author and Staden further developed the absorption method for graph collections, combined with the application of the transversal blow-up lemma, they finally proved a stability version of Joos and Kim’s result [21]. The absorption technique used in [10] originates from [12] and [8], based on the common absorption technique introduced first in [25], which was showed to be extremely powerful in the solutions of various transversal graph embedding problems. In this paper, we adapt the method developed in [10] and generalize their stability from graph collections to digraph collections. By analyzing the extremal cases, we proved a tight bound for the transversal Ghouila-Houri’s theorem. We believe our method could be used to demonstrate stabilities or tight bounds for other transversal graph embedding problem (see the survey [26]).
@@
Our proof utilizes the regularity blow-up method in digraph collections, and is divided into two parts: stable case and the extremal case.
Stable case. We first characterize the extremality of a single digraph, which forms the basis of our extremal case distinction. Let $0<\frac{1}{n}\ll\epsilon\ll 1$ and $D$ be an $n$-vertex digraph with minimum semi-degree at least $(\frac{1}{2}-\epsilon)n$. If $D$ contains no directed Hamilton cycle, then $D$ contains a sub-digraph that is close to one of the digraphs in Figure 1. Based on Lemma 3.3, there exists a characteristic partition of $D$ in each case, and we say that $D$ is $(\epsilon,{\rm EC}k)$-extremal for some $k\in[3]$. Moreover, the digraph $D$ can be $(\epsilon,{\rm EC}k)$-extremal for exactly one $k$.
[…]
@@
We build an “directed absorbing rainbow cycle” $C$ for $\mathcal{D}$ with the property that $C$ is very small and there is a set $A$ consisting of some color-vertex-vertex triples $(c,u,v)$ such that whenever $A_{0}\subseteq A$ is sufficiently small compared to $|C|$, $C$ can absorb all of its elements. Additionally, we construct an additional auxiliary set $\mathcal{Q}$ consisting of color-color-vertex-vertex tuples $(c,c^{\prime},v,v^{\prime})$ to absorb those colors and vertices which $C$ cannot. Delete all colors and vertices in $C$ and $\mathcal{Q}$ from $\mathcal{D}$.
@@
Step 2. Use the regularity-blow-up method for transversals in digraph collections to cover almost all vertices and colors with long directed paths.
@@
We first prove the regularity lemma for digraph collections (Section 2.1), and then apply it to $\mathcal{D}$, yielding a reduced digraph collection $\mathcal{R}$ that inherits the minimum degree condition and stability of $\mathcal{D}$. By considering the characteristic bipartite graph of each digraph, we get a bipartite graph collection $\mathcal{B_{R}}$ for the reduced digraph collection $\mathcal{R}$. Notice that $\mathcal{R}$ has a disjoint rainbow directed cycle partition if and only if $\mathcal{B_{R}}$ has a transversal perfect matching. We establish a stability result for transversal perfect matching in bipartite graph collections (Section 2.2), ensuring that the disjoint rainbow directed cycle partition of $\mathcal{R}$ exists. Using the blow-up method, we obtain a set of almost spanning disjoint rainbow directed paths within each directed cycle of $\mathcal{R}$, which covers almost all vertices outside $C$ and $\mathcal{Q}$ (Section 4).
@@
Step 3. Connect the obtained directed paths and cover remaining vertices via the absorbing cycle.
@@
In this step, we apply the absorbing property of $C$ as well as the auxiliary set $\mathcal{Q}$ to connect all the directed paths and cover the remaining vertices, ultimately form a transversal directed Hamilton cycle.
[…]
Firstly, expand $A_{1}\cup B_{1}$ into an equitable partition $A\cup B$ of $V$. For a vertex $v\in A$, there are three bad cases (similar for vertices in $B$): $v$ lies in almost all $B_{i}$ for $i\in\mathcal{C}_{1}\cup\mathcal{C}_{2}$; $v$ lies in almost all $L_{i}$ for $i\in\mathcal{C}_{1}\cup\mathcal{C}_{2}$; $v\in A_{i}$ for many $i\in\mathcal{C}_{1}\cup\mathcal{C}_{2}$ and $v\in B_{j}$ for many $j\in\mathcal{C}_{1}\cup\mathcal{C}_{2}$. Move vertices in the first case from $A$ to $B$, and do the same operation for such vertices in $B$. Denote the set of vertices in the other two bad cases by $V_{\rm bad}$. After this, the number of vertices in $A$ and $B$ may be unbalanced, but their difference is still small. Without loss of generality, assume that $|A|\leq|B|$. In order to apply the transversal blow-up lemma to find long rainbow directed paths inside $\mathcal{D}$, the following four steps are required.
@@
If there exists a set of disjoint rainbow paths in $\mathcal{D}[B]$ such that after deleting their vertices, the number of vertices in $A$ and $B$ is balanced, then we are done. Otherwise, we will show that $\mathcal{D}$ must contains a transversal directed Hamilton cycle (see Lemma 5.14).
@@
Using colors in $\mathcal{C}_{2}$, we are to find a series of disjoint rainbow directed paths $P_{3}$ such that their centers are all vertices in $V_{\rm bad}$ and endpoints are unused vertices in $(A\cup B)\setminus V_{\rm bad}$.

### v2

@@ 1 Introduction
A transversal $X$ over a collection $\mathcal{F}=\{F_{1},\ldots,F_{m}\}$ of objects represents an object that intersects every $F_{i}$. Here, the object can be any mathematical object, such as sets, spaces, set systems, matroids, and so forth. Several classical theorems, which guarantee the existence of certain objects, have been restated in the context of transversals, illustrating the feasibility or infeasibility of obtaining such an object as a transversal over a certain collection. To name a few, transversals are extensively studied for Carathéodory’s theorem [4], Helly’s theorem [22], Erdős-Ko-Rado theorem [2], Rota’s basis conjecture [20, 24], and others.
@@
For a given collection $\mathcal{G}=\{G_{1},\ldots,G_{m}\}$ of graphs/hypergraphs/digraphs with the same vertex set $V$, an $m$-edge graph/hypergraph/digraph $H$ with vertices in $V$ is transversal in $\mathcal{G}$ if there exists a bijection ${\varphi}:E(H)\to[m]$ such that $e\in E(G_{{\varphi}(e)})$ for all $e\in E(H).$
By interpreting each $G_{i}$ as the set of edges colored with color $i$, the function $\varphi$} is often referred to as a coloring. We say that the coloring $\varphi$} is rainbow if it is injective, and $H$ is a rainbow subgraph if there exists a rainbow coloring on it. Many classical results in extremal graph theory have been extended to such transversal settings, exhibiting interesting phenomena. The central question in studying transversal generalizations can be stated as follows.
@@
Let $H$ be a graph/hypergraph/digraph with $m$ edges, $\mathcal{G}=\{G_{1},G_{2},\dots,G_{m}\}$ be a collection of graphs/hypergraphs/digraphs on a common vertex set of size $n$. Which property $\mathcal{P}$ imposed on $\mathcal{G}$ will guarantee the existence of a transversal copy of $H$?
Since every object in the family $\mathcal{G}$ can be the same, a necessary condition for a positive answer to the above question is that every $n$-vertex graph/hypergraph/digraph satisfying the property $\mathcal{P}$ contains a copy of $H$. However, this alone is not always sufficient. For example, Aharoni, DeVos, de la Maza, Montejano, and Šámal [1] proved that if $\mathcal{G}=\{G_{1},G_{2},G_{3}\}$ is a collection of graphs on a common vertex set of size $n$ and $|E(G_{i})|>(\frac{26-2\sqrt{7}}{81})n^{2}$ for all $i\in[3]$, then $\mathcal{G}$ contains a rainbow triangle. Moreover, the constant $\frac{26-2\sqrt{7}}{81}$ is optimal, while Mantel’s theorem states that any $n$-vertex graph with more than $\lfloor\frac{n^{2}}{4}\rfloor$ edges must contain a triangle.
@@
A natural next step is to study the rainbow analogues in digraph collections. Cheng, Han, Wang and Wang [8] studied the transversal version of Hajnal-Szemerédi theorem in digraph collections, which asymptotically generalized the corresponding results by Czygrinow-DeBiasio-Kierstead-Molla [15] and Treglown [30]. It is well-known that every tournament contains a Hamilton path, and every strongly connected tournament contains a Hamilton cycle. Chakraborti, Kim, Lee and Seo [6] proved the existence of a transversal directed Hamilton path and cycle in a family of tournaments. They [7] also proved that in a digraph collection consisting of tournaments, there exist transversal Hamilton cycles of all possible orientations possibly except the consistently oriented one. These results generalized the classical theorem of Thomason [28]. Babiński, Grzesik and Prorok [3] studied the transversal Mantel’s theorem in digraph and oriented graph collections. Gerbner, Grzesik, Palmer and Prorok [18] determined the minimum number of edges in each color which guarantees the existence of any rainbow directed star.
@@ 1.1 Main result
In this paper, we study the transversal problem in digraph collections. Let $D$ be a digraph with no loops and at most one edge in each direction between every pair of vertices. Denote the set of vertices of $D$ by $V(D)$. Define $\delta^{+}(D)$ and $\delta^{-}(D)$ as the minimum out-degree and in-degree of $D$, respectively, and let the minimum semi-degree be $\delta^{0}(D)=\min{\delta^{+}(D),\delta^{-}(D)}$. For a collection $\mathcal{D}=\{D_{1},\dots,D_{m}\}$ of digraphs (not necessarily distinct) on a common vertex set $V$, define $\delta^{0}(\mathcal{D})=\min\{\delta^{0}(D_{i}):i\in[m]\}$.
In 1960, Ghouila-Houri [19] generalized Dirac’s theorem to digraphs.
@@
For each integer $n\geq 3$, every $n$-vertex digraph $D$ with $\delta^{0}(D)\geq\frac{n}{2}$ contains a directed Hamilton cycle.
Chakraborti, Kim, Lee, and Seo [6] suggested that it would be interesting to consider a transversal version of Ghouila-Houri’s theorem. Our main result solves this.
@@
In an undirected graph, one may replace each undirected edge by two arcs with opposite directions. Thus, for sufficiently large $n$, Theorem 1.2 is a direct consequence of Theorem 1.4. Furthermore, by setting $D_{1}=D_{2}=\cdots=D_{n}$ in Theorem 1.4, we immediately obtain Theorem 1.3 for all sufficiently large $n$.
@@ 1.2 Notation and Organization
We use $H\subseteq D$ to denote that $H$ is a sub-digraph of $D$. For $S\subseteq V(D)$, we write $D[S]$ for the sub-digraph of $D$ induced by $S$ and DSD-S for $D[V(D)\setminus S]$. For two not necessarily disjoint vertex subsets $A,B\subseteq V(D)$, let $D[A,B]$ be the sub-digraph of $D$ with vertex set $A\cup B$ and edge set $\{xy\in E(D):x\in A,y\in B\}$ and let $e_{D}(A,B):=|E(D[A,B])|$. Denote $D^{\pm}[A,B]=D[A,B]\cup D[B,A]$. When $A=B$, we often abbreviate $D[A,B]$ and $e_{D}(A,B)$ as $D[A]$ and $e_{D}(A)$ respectively.
For a digraph $D$ and $v\in V(D)$, we denote the sets of out-neighbors and in-neighbors of $v$ by $N_{D}^{+}(v)$ and $N_{D}^{-}(v)$, respectively, with the corresponding out- and in-degrees given by $d_{D}^{+}(v):=|N_{D}^{+}(v)|$ and $d_{D}^{-}(v):=|N_{D}^{-}(v)|$. For $S\subseteq V(D)$, we write $d_{D}^{+}(v,S):=|N_{D}^{+}(v)\cap S|$ and $d_{D}^{-}(v,S):=|N_{D}^{-}(v)\cap S|$. We often omit subscripts when they are clear from the context.
Let $\mathcal{D}=\{D_{i}:i\in[m]\}$ be a collection of digraphs on a common vertex set $V$ and $H$ be a rainbow digraph inside $\mathcal{D}$. For a vertex $v\in V$ and each $i\in[m]$, let $N_{i}^{+}(v):=N_{D_{i}}^{+}(v)$, $N_{i}^{-}(v):=N_{D_{i}}^{-}(v)$, $d_{i}^{+}(v):=|N_{i}^{+}(v)|$ and $d_{i}^{-}(v):=|N_{i}^{-}(v)|$. For any two vertex subsets $X,\,Y\subseteq V$, define $\mathcal{D}[X]:=\{D_{i}[X]:i\in[m]\}$, $\mathcal{D}[X,Y]:=\{D_{i}[X,Y]:i\in[m]\}$ and $\mathcal{D}^{\pm}[X,Y]:=\{D_{i}^{\pm}[X,Y]:i\in[m]\}$. If $E(D)=\emptyset$, then we simply write $D=\emptyset$. Similarly, if there is no edge in $D_{i}$ for all $i\in[m]$, then we write $\mathcal{D}=\emptyset$. Denote by ${\rm col}(H)$ the set of colors appearing in $H$. Two rainbow graphs $H_{1}$ and $H_{2}$ are said to be disjoint if $V(H_{1})\cap V(H_{2})=\emptyset$ and ${\rm col}(H_{1})\cap{\rm col}(H_{2})=\emptyset$.
@@
Organization. We conclude this section with a sketch of the proof of Theorem 1.4. In Section 2, we introduce two key techniques that we will need later, i.e., the regularity lemma in digraph collections and a stability result for transversal perfect matchings in bipartite graph collections. For readability, we postpone their proofs to the appendix. The proof of Theorem 1.4 will be divided into “stable case” and “extremal case”. Section 3 introduces notions of extremality and stability for digraphs and digraph collections; we prove several results about absorption that will be used in the proof of the stable case. In Section 4, we show that there is always a transversal directed Hamilton cycle in the stable case. Sections 5-7 deal with the extremal case. We finish our paper with some concluding remarks in the final section.
@@ 1.3 Discussion of the strategy
Joos and Kim [21] proved the transversal version of Dirac’s theorem by introducing the auxiliary digraph technique, along with some rotations. However, this technique is not feasible in digraph collections because we cannot extend an existing directed cycle into a longer cycle by using the same rotation. In [9], the first author and Staden further developed the absorption method for graph collections, combined with the application of the transversal blow-up lemma, they finally proved a stability version of Joos and Kim’s result [21]. The absorption technique in [9] originates from [12] and [8], which adapted the absorption technique first introduced in [25]. This approach has become a powerful tool for solving various transversal graph embedding problems.
In this paper, we adapt the method developed in [9] and generalize their stability from graph collections to digraph collections. By analyzing the extremal cases, we proved a tight bound for the transversal Ghouila-Houri’s theorem. We believe that our method could be used to demonstrate stabilities or tight bounds for other transversal graph embedding problems (see the survey [26]).
@@
Our proof utilizes the regularity blow-up method in digraph collections, and is divided into two parts: stable case and extremal case.
Stable case. We first characterize the extremality of a single digraph, which forms the basis of our extremal case distinction. Let $0<\frac{1}{n}\ll\epsilon\ll 1$ and $D$ be an $n$-vertex digraph with minimum semi-degree at least $(\frac{1}{2}-\epsilon)n$. Suppose $D$ has no directed Hamilton cycle. Then $D$ contains a sub-digraph that is close to one of the digraphs in Figure 1. By Lemma 3.2, such a sub-digraph induces a characteristic partition of $D$; we say that $D$ is $(\epsilon,\mathrm{EC}k)$-extremal for some $k\in[3]$. Furthermore, this $k$ is unique.
Assume that $\mathcal{D}=\{D_{1},\ldots,D_{n}\}$ is a collection of digraphs with the same vertex set $V$ of size $n$. Suppose that $\delta^{0}(\mathcal{D})\geq(\frac{1}{2}-\epsilon)n$ and $\mathcal{D}$ contains no transversal directed Hamilton cycles. We suspect that either almost all digraphs in $\mathcal{D}$ are $(\epsilon,{\rm EC}k)$-extremal for some $k\in[2]$ (allowing $k=1$ for some digraphs and $k=2$ for others), or almost all digraphs in $\mathcal{D}$ are $(\epsilon,{\rm EC}3)$-extremal; moreover those digraphs have similar vertex partitions. Based on this observation, we define two kinds of stability for a digraph collection with minimum degree at least $(\frac{1}{2}-\epsilon)n$. We say that
@@
We build a “directed absorbing rainbow cycle” $C$ for $\mathcal{D}$ with the property that $C$ is very small and there is a set $A$ consisting of some color-vertex-vertex triples $(c,u,v)$ such that whenever $A_{0}\subseteq A$ is sufficiently small compared to $|C|$, $C$ can absorb all of its elements. Additionally, we construct an additional auxiliary set $\mathcal{Q}$ consisting of color-color-vertex-vertex tuples $(c,c^{\prime},v,v^{\prime})$ to absorb those colors and vertices which $C$ cannot. Delete all colors and vertices in $C$ and $\mathcal{Q}$ from $\mathcal{D}$.
@@
Step 2. Use the regularity-blow-up method for transversals in digraph collections to cover almost all vertices and colors with long rainbow directed paths.
@@
We first prove a regularity lemma for digraph collections (Section 2.1), and then apply it to $\mathcal{D}$, yielding a reduced digraph collection $\mathcal{R}$ that inherits the minimum degree condition and the stability property of $\mathcal{D}$. By considering the characteristic bipartite graph of each digraph, we get a bipartite graph collection $\mathcal{B_{R}}$ for the reduced digraph collection $\mathcal{R}$. Notice that $\mathcal{R}$ has a disjoint rainbow directed cycle partition if and only if $\mathcal{B_{R}}$ has a transversal perfect matching. We establish a stability result for transversal perfect matching in bipartite graph collections (Section 2.2), ensuring that the disjoint rainbow directed cycle partition of $\mathcal{R}$ exists. Using the blow-up method, we obtain a set of almost spanning disjoint rainbow directed paths within each directed cycle of $\mathcal{R}$, which covers almost all vertices outside the absorbing cycle $C$ and the auxiliary set $\mathcal{Q}$ (Section 4).
@@
Step 3. Connect the obtained directed paths and cover the remaining vertices via the absorbing cycle.
@@
In this step, we apply the absorbing property of $C$ as well as the auxiliary set $\mathcal{Q}$ to connect all the directed rainbow paths and to cover the remaining vertices, which ultimately forms a transversal directed Hamilton cycle.
Extremal case (Sections 5-7). Assume that $\delta^{0}(\mathcal{D})\geq\frac{n}{2}$ and $\mathcal{D}$ contains no transversal Hamilton cycles. Let $0<\frac{1}{n}\ll\epsilon\ll\delta\ll\eta\ll 1$. Denote $\mathcal{C}_{k}:=\{i\in[n]:D_{i}\ {\text{is ($\epsilon$,EC$k$)-extremal}}\}$ for $k\in[3]$ and $\mathcal{C}_{\rm bad}:=[n]\setminus(\mathcal{C}_{1}\cup\mathcal{C}_{2}\cup\mathcal{C}_{3})$. Notice that $\mathcal{C}_{i}\cap\mathcal{C}_{j}=\emptyset$ for $1\leq i<j\leq 3$. By “stable case”, we know that almost all digraphs in $\mathcal{D}$ are extremal and $|\mathcal{C}_{1}|+|\mathcal{C}_{2}|\leq 2\sqrt{\delta}n$ if $|\mathcal{C}_{3}|\geq 2\sqrt{\delta}n$. In view of Lemma 3.2, we can fix a characteristic partition $({A_{i}},{B_{i}},{L_{i}})$ for each graph $D_{i}$ with $i\in\mathcal{C}_{1}\cup\mathcal{C}_{2}$, and a characteristic partition $({C_{i}^{1}},C_{i}^{2},C_{i}^{3},C_{i}^{4},L_{i})$ for each graph $D_{i}$ with $i\in\mathcal{C}_{3}$. By swapping labels, we know that $|{A_{1}}\triangle{A_{i}}|,|{B_{1}}\triangle{B_{i}}|<\delta n$ for every $i\in(\mathcal{C}_{1}\cup\mathcal{C}_{2})\setminus\{1\}$ and $|C_{1}^{k}\triangle C_{i}^{k}|<2\delta n$ for every $i\in\mathcal{C}_{3}\setminus\{1\}$. We proceed with our proof by considering the sizes of $|\mathcal{C}_{1}|$, $|\mathcal{C}_{2}|$ and $|\mathcal{C}_{3}|$. We will take the case $|\mathcal{C}_{1}|<\eta n$, $|\mathcal{C}_{2}|\geq\eta n$ and $|\mathcal{C}_{3}|\leq 2\sqrt{\delta}n$ as an example to illustrate the general idea of our proof.
Firstly, expand $A_{1}\cup B_{1}$ into an equitable partition $A\cup B$ of $V$. To apply the transversal blow-up lemma (Lemma 6.2) for finding long rainbow directed paths in $\mathcal{D}$, we need to ensure most vertices in the partition $A\cup B$ satisfy uniform neighborhood properties (for example, a vertex $v\in A$ should satisfy $v\in A_{i}$ for the majority of colors $i\in\mathcal{C}_{2}$). We identify three vertex types in $A$ (and symmetrically in $B$) that may disrupt the desired uniformity:
  1. Type 1.
    $v\in A$ belongs to $L_{i}$ for many $i\in\mathcal{C}_{2}$;
  2. Type 2.
    $v\in A$ belongs to $A_{i}$ for many $i\in\mathcal{C}_{2}$, and $v\in B_{j}$ for many other $j\in\mathcal{C}_{2}$;
  3. Type 3.
    $v\in A$ belongs to nearly all $B_{i}$ for $i\in\mathcal{C}_{2}$.
We handle these cases as follows: For Type 3 vertices in $A$, we move them to $B$ (and perform the analogous operation for Type 3 vertices in $B$). We define $V_{\rm bad}$ as the union of vertices of Type 1 and Type 2. After this reallocation, every vertex in $A\setminus V_{\rm bad}$ (resp. $B\setminus V_{\rm bad}$) lies in $A_{i}$ (resp. $B_{i}$) for the majority of colors $i\in\mathcal{C}_{2}$. However, the partition sizes $|A|$ and $|B|$ may become unbalanced, with their difference bounded by $9\delta n$. To balance them, we require the following four steps.
@@
If there exists a set of disjoint rainbow paths in $\mathcal{D}[B]$ such that after deleting their vertices, the number of vertices in $A$ and $B$ is balanced, then we are done. Otherwise, we will show that $\mathcal{D}$ must contain a transversal directed Hamilton cycle (see Lemma 6.5).
@@
Using colors in $\mathcal{C}_{2}$, we are to find a sequence of disjoint rainbow directed paths $P_{3}$ such that their centers are all vertices in $V_{\rm bad}$ and endpoints are unused vertices in $(A\cup B)\setminus V_{\rm bad}$.